OpenStudy (anonymous):

how to integrate (x^2-1)/(x+1)?

OpenStudy (saifoo.khan):

\[\int\limits \frac{x^2-1}{x+1} \to \int\limits \frac{(x-1)(x+1)}{x+1}\]

OpenStudy (tkhunny):

Of course, you DO need to stay away from \(x = -1\).